# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@ 9.4 Challenges
Many existing methods rely on prompts to generate user-friendly explanations, which are not as reliable as numerical methods with good theoretical foundations. Ye & Durrett (2022b) find that the explanations by CoT may not be factually grounded in the inputs. Therefore, they believe that these explanations are more suitable as post-hoc explanations regarding why the LLM makes certain predictions (no matter whether the predictions are right or wrong). However, the validity of viewing CoT explanations as post-hoc justifications has been questioned by recent findings from Turpin et al. (2024), which uses biased datasets (e.g., the few-shot examples in the prompt always answer “A” for multiple choice questions) to show that the generated explanations may be plausible, but systematically unfaithful to represent the true reasoning process of the LLMs. This issue arises from biased reasoning steps which inadvertently skew the predictions. Thus, there’s a growing need for more theoretical scrutiny of user-friendly explanations to ensure faithfulness and credibility.
9.4.2 Constrained Application Scenarios
@@
Currently, the utilization of LLMs to explain smaller black-box models is mainly limited to those that deal with data with rich textual information (Bhattacharjee et al., 2023; Lei et al., 2023). Although Zhao et al. (2023b) propose a strategy to explain image classifiers, the ability to match candidate textual explanations with image patterns still relies on the pretrained vision-language model CLIP. This method may not be applicable to other domains such as graph machine learning (e.g., graph neural networks), or time-series analysis (recurrent neural networks), where large pretrained models have demonstrated little progress compared to natural language processing and computer vision. Therefore, there is a compelling need to devise more versatile strategies for explaining models across a wider range of fields. This endeavor could depend on the fundamental research on combining LLM with other domain-specific tasks, such as the development of Graph-Language Models that are applicable to unseen graphs in a zero-shot manner.
10 LLMs for Interpretable AI System Design
@@
An intriguing but challenging problem in XAI is creating model architectures or even AI systems that are inherently interpretable (Rudin, 2019), where different model components represent clear and comprehensible concepts or functionalities that are easily distinguishable from one another. Machine learning models such as support vector machines (Hearst et al., 1998) and tree-based models (Song & Ying, 2015) were classical techniques for achieving model interpretability. In the deep learning era, typical research areas in this context include concept-bottleneck models (Koh et al., 2020; Yuksekgonul et al., 2022), disentangled representation learning (Denton et al., 2017; Higgins et al., 2016), and network dissection (Bau et al., 2017; 2018). Nevertheless, under the traditional deep learning setting, the usability of these techniques remains limited because of two major challenges. First, it is difficult to define the spectrum of concepts or functionalities the model is expected to capture. Second, the efficacy of interpretable models often falls short compared to black-box models, thereby constraining their practical utility.
@@
Large foundation models, such as large language models (LLMs) and vision language models (VLMs), provide opportunities to bridge the gap. By leveraging the common-sense knowledge embedded within them, foundation models can design interpretable architectures by providing cues that encourage creating and using the features or procedures within AI workflows. This is different from traditional deep learning pipelines where the deep models automatically discover the features during the training process, which may not end up with model components with clear meanings. Furthermore, LLMs can decompose complex tasks into simpler and collaborative sub-tasks, enhancing both the system’s interpretability and its overall performance.
10.1 Designing Interpretable Network Architectures with LLMs
@@
Representative methods for developing interpretable deep architectures include Generalized Additive Models (GAMs) (Zhuang et al., 2021; Lou et al., 2012) and Concept Bottleneck Models (CBMs) (Koh et al., 2020; Yuksekgonul et al., 2022). These models map inputs into a human-understandable latent space, and then apply a linear transformation from this space to the target label. For example, to build a classifier that diagnoses arthritis, we can let the model identify features such as “bone spurs” and “sclerosis”, and then use these interpretable features for the final decision. However, these approaches often require the involvement of experts to define the latent space, which can limit the learning capabilities of deep models. Some work tries to automate the discovery of semantic concepts during model training, such as by requiring independence between concepts (Higgins et al., 2016; Yu et al., 2020) or clustering data (Ghorbani et al., 2019), but they lack direct control over the outcomes and does not ensure the clarity of the concepts. One promising strategy is to utilize LLMs to provide comprehensible concept candidates. Menon & Vondrick (2022) use human language as an internal representation for visual recognition, and create an interpretable concept bottleneck for downstream tasks. By basing the decision on those comprehensible concepts, the model architecture itself is provided with better transparency. Similarly, a recent approach Labo (Yang et al., 2023b) constructs high-performance CBMs without manual concept annotations. This method controls the concept selection in bottlenecks by generating candidates from the LLMs, which contain significant world knowledge (Petroni et al., 2019) that can be explored by prompting a string prefix. Human studies further indicate that those LLM-sourced bottlenecks are much factual and groundable, maintaining great inherent interpretability for model designs. Besides the concept-based models, another promising strategy is to employ LLMs to enhance the conventional architectures that are inherently interpretable, such as GAMs and Decision Trees (DTs). Singh et al. (2023) leverages the knowledge captured in LLMs to enhance GAMs and DTs, where LLMs are only involved during the augmented model training instead of the inference process. For GAMs training, LLMs can provide decoupled embeddings for enhancement. For DTs training, LLMs are able to help generate improved features for splitting. The LLM-augmented GAMs and DTs enable full transparency, where only the summing coefficients and input key phrases are required for interpretation. With the extra information from LLMs, augmented GAMs and DTs are capable of achieving better generalization performance compared with non-augmented ones.
10.2 Designing Interpretable AI Workflows with LLM Agents
@@
Traditional deep models are usually designed in an end-to-end manner. The internal workflows are not quite understandable to general users. By utilizing common-sense world knowledge, LLMs can break down complex problems into smaller ones and organize the workflows among them, leading to more interpretable design of AI systems (Feng et al., 2023). A recent example on interpretable AI workflow design comes from  Shen et al. (2024), where an LLM-powered agent leverages ChatGPT to integrate various off-the-shelf AI models (e.g., from Hugging Face (Jain, 2022)) to handle different downstream application tasks. In order to handle the complicated tasks in a transparent workflow, LLMs serve as a pivotal role in coordinating with external models with language mediums to harness their powers. By planning the target task, selecting candidate models, executing decomposed subtasks and summarizing responses, LLMs can help disassemble tasks based on user requests, and assign appropriate models to the tasks based on the model descriptions. Similarly, to transparentize the workflow, Liu et al. (2023d) introduces a task decomposer to analyze the user prompts and break it down into a number of subtasks for solving using LLMs. Each subtask is well managed and attributed with description, domain, inputs, and outputs. In this way, the AI systems are then capable of handling intricate user prompts with a step-by-step understandable workflow. Under the prompting paradigm, Khot et al. (2022) also employs LLMs to solve complex tasks by decomposition. Drawing inspiration from software libraries where the workflows are trackable, the decomposer and shared subtasks are designed in a modular manner. One step further, Wang et al. (2024) introduces an interactive planning approach for complex tasks, which enhances the error correction on initial LLM-generated plans by integrating plan execution descriptions and providing self-explanation of the feedback. Such interactive nature enables better workflow transparency in long-term planning and multi-step reasoning task scenarios.
10.3 Challenges
10.3.1 Planning Feasibility in Complicated Scenarios
@@
Despite the task planning capability of LLMs, it is still challenging to be applied to certain scenarios in real-world applications due to the feasibility issues. One typical scenario is the few-shot planning cases (Guo et al., 2023), where acquiring large datasets for training is either impractical or cost-prohibitive thus making feasible planning about unseen cases from sparse exemplars extremely challenging. To better assist the interpretable designs, LLM planning needs to generalize well without extensive supervision and is expected to have the ability to integrate information from prior experiences as well as knowledge. Besides, another important scenario lies in the dynamic planning settings (Dagan et al., 2023), in which LLMs integrate feedback from the environment iteratively, letting the agent take thinking steps or augment its context with a reasoning trace. Dynamic scenarios urgently and frequently involve high computational costs resulting from the iterated invocations of LLMs, and still face challenges in dealing with the limits of the context window and recovering from hallucinations on planning.
10.3.2 Assistance Reliability with Knowledge Gaps
@@
LLMs exhibit remarkable proficiency in encapsulating real-world knowledge within their parameters, but they resort to hallucinations and biases with high confidence when certain knowledge is missing or unreliable. Although a growing number of techniques has been proposed, such as retrieval augmentation (Guu et al., 2020), searching integration (Nakano et al., 2021) and multi-LLM collaboration (Feng et al., 2023), to expand LLM knowledge, such discrepancy in knowledge may perpetually exist owing to the continuously evolving character of human understanding (Ji et al., 2023). As a result, a crucial research challenge keeps rising, i.e., how to effectively detect and mitigate the LLM knowledge gaps from humans when employing LLMs for designs. We will need further research on evaluating and developing robust LLM mechanisms to address the knowledge-gapping problems, with the goal of helping improve LLM reliability, reducing hallucinations and mitigating biases. Furthermore, the intersections between the knowledge gaps and the safety aspects are also of great challenges to be solved, which may pose some security concerns especially when using LLMs for downstream models or workflow designs.
11 Emulating Humans with LLMs for XAI
@@
This section discusses how LLMs can be leveraged to serve XAI by playing the role of humans. Building explainable models requires two main steps where humans are in the loop: (1) collecting a dataset with human-annotated rationales to train the models; (2) collecting human feedback on the quality of explanations produced by the models for evaluation. The significant cost and time required for human involvement raise the main challenge in scaling up this procedure. LLMs emerge as a promising solution to this challenge, thanks to their capability to emulate human reasoning and produce responses that closely resemble human-generated content. In the following, we introduce the methods that demonstrate LLMs’ ability to generate human-like annotations and feedback, contributing to the creation of explainable models.
11.1 Emulating Human Annotators for Training Explainable Models
@@
Incorporating human-understandable rationales into model development has shown its effectiveness in enhancing both the transparency and performance of the system for various NLP tasks, such as question answering (Li et al., 2018; Wu et al., 2020a), sentiment analysis (Du et al., 2019b; Antognini & Faltings, 2021), and common sense reasoning (Rajani et al., 2019; Camburu et al., 2021). We use the term rationales to describe supportive evidence that justifies the connection between inputs and outputs (Gurrapu et al., 2023). Traditionally, the rationales are collected by leveraging human annotations (Camburu et al., 2018; Wang et al., 2019) or applying expert-designed rules (Alhindi et al., 2018; Li et al., 2018), resulting in expensive costs or limited quality. Recently, researchers in automatic annotation (Ding et al., 2022; Belal et al., 2023; Gilardi et al., 2023) have begun to explore the potential of leveraging advanced LLMs to emulate human annotators in annotating the target labels of task-specific examples. These studies found that advanced LLMs show comparable annotation qualities against average crowd human annotators on most tasks with a lower cost, pointing out the scalability of using machine-emulated annotators. Inspired by these works, some studies (Huang et al., 2023b; a) attempt to leverage advanced LLMs to collect rationales by applying the chain-of-thought technique. Specifically, researchers provide several input-rationale-output demonstrations within the input text to prompt the LLMs to generate rationale and output for an unlabeled input instance. The quality of such annotated rationales largely relies on the in-context learning capabilities of LLMs, leading to uncontrollable annotation quality on uncommon tasks. Other scholars (Yao et al., 2023a; Chen et al., 2023b; Luo et al., 2023) propose a human-in-the-loop LLM-based annotation framework based on the active-learning architecture. This framework initially collects a small seed dataset with human-annotated rationales and labels. This seed dataset is used to train an explainable classifier for this downstream task. Then, each unlabeled sample is passed through the trained explainable classifier. This is followed by a selection strategy that chooses representative samples according to metrics such as explanation plausibility, prediction uncertainty, and sample diversity. Finally, LLMs are leveraged to annotate the rationales and labels of these selected unlabeled samples. This procedure could be repeated multiple times, and the trained explainable classifier from the latest time is the final output of this framework. Compared with other methods, this approach balances the annotation quality and the cost budget in developing explainable models by using LLM-emulated annotators.
11.2 Emulating Human Feedback for Evaluating Explainable Models
@@
The explanations generated by the explainable models could be classified into two categories: extractive and abstractive (Gurrapu et al., 2023). Extractive explanations derive directly from the input data, exemplified by attribution-based methods that emphasize specific segments of the input text. In contrast, abstractive explanations are generated in a free-form text manner, such as chain-of-thought (CoT) responses (Wei et al., 2022), offering a more nuanced interpretation. The quality of extractive explanations is typically assessed through their agreement with annotated rationales (DeYoung et al., 2020), such as accuracy, recall, and precision. However, evaluating abstractive explanations presents a significant challenge, as it is impractical to exhaustive all reasonable abstractive results comprehensively. To automatically assess abstractive explanations, early studies first collect some free-text rationales, and then apply LLMs to estimate the similarity between the explanation and the rationales (Cheng et al., 2023; Li et al., 2023b). A higher similarity between the abstraction explanation and the annotated rationales indicates a more transparent model. Recently, some researchers directly use LLMs to check the rationality of the model explanations without referring to human-annotated rationales (Miao et al., 2023; Bills et al., 2023), emphasizing the potential of emulating human feedback with advanced LLMs.
11.3 Challenges
11.3.1 Uncontrollable Credibility of Emulation

### v2

@@ 9.3. Challenges
Many existing methods rely on prompts to generate user-friendly explanations, which are not as reliable as numerical methods with good theoretical foundations. (Ye and Durrett, 2022) find that the explanations by CoT may not be factually grounded in the inputs. Therefore, they believe that these explanations are more suitable as post-hoc explanations regarding why the LLM makes certain predictions. However, the validity of viewing CoT explanations as post-hoc justifications has been questioned by recent findings (Turpin et al., 2024), which uses biased datasets (e.g., the few-shot examples in the prompt always answer “A” for multiple choice questions) to show that the generated explanations may be plausible, but systematically unfaithful to represent the true reasoning process of the LLMs. Therefore, there’s a growing need for more theoretical scrutiny of user-friendly explanations to ensure faithfulness and credibility.
9.3.2. Constrained Application Scenarios
@@
Currently, the utilization of LLMs to explain smaller black-box models is mainly limited to those that deal with data with rich textual information (Bhattacharjee et al., 2023; Lei et al., 2023). Although (Zhao et al., 2023a) propose a strategy to explain image classifiers, the ability to match candidate textual explanations with image patterns still relies on the pretrained vision-language model CLIP. Therefore, there is a compelling need to development more versatile strategies for explaining models across a wider range of fields. This endeavor could depend on the fundamental research on combining LLM with other domain-specific tasks, such as the development of Graph-Language Models that are applicable to unseen graphs in a zero-shot manner.
10. Interpretable AI System Design with LLMs
@@
An intriguing but challenging problem in XAI is creating model architectures or even AI systems that are inherently interpretable (Rudin, 2019), where different model components represent clear and comprehensible concepts or functionalities that are easily distinguishable from one another. Machine learning models such as support vector machines (Hearst et al., 1998) and tree-based models (Song and Ying, 2015) were classical techniques for achieving model interpretability. In the deep learning era, typical research areas in this context include concept-bottleneck models (Koh et al., 2020; Yuksekgonul et al., 2022), disentangled representation learning (Denton et al., 2017; Higgins et al., 2016), and network dissection (Bau et al., 2017, 2018). Nevertheless, under the traditional deep learning setting, the usability of these techniques remains limited because of two major challenges. First, it is difficult to define the spectrum of concepts or functionalities the model is expected to capture. Second, the efficacy of interpretable models often falls short compared to black-box models, thereby constraining their practical utility.
@@
Large foundation models, such as large language models (LLMs) and vision language models (VLMs), provide opportunities to bridge the gap. By leveraging the common-sense knowledge embedded within them, foundation models can design interpretable architectures by providing cues that encourage creating and using the features or procedures within AI workflows. This is different from traditional deep learning pipelines, where the deep models automatically discover the features during the training process, which may not end up with model components with clear meanings. Furthermore, LLMs can decompose complex tasks into simpler and collaborative sub-tasks, enhancing both the system’s interpretability and its overall performance.
10.1. Designing Interpretable Network Architectures with LLMs
@@
Representative methods for developing interpretable deep architectures include Generalized Additive Models (GAMs) (Agarwal et al., 2021; Zhuang et al., 2021; Lou et al., 2012) and Concept Bottleneck Models (CBMs) (Koh et al., 2020; Yuksekgonul et al., 2022; Shang et al., 2024). These models map inputs into a human-understandable latent space, and then apply a linear transformation from this space to the target label. For example, to build a classifier that diagnoses arthritis, we can let the model identify features such as “bone spurs” and “sclerosis”, and then use these interpretable features for the final decision. However, these approaches often require the involvement of experts to define the latent space, which can limit the learning capabilities of deep models. Some work tries to automate the discovery of semantic concepts during model training, such as by requiring independence between concepts (Higgins et al., 2016; Yu et al., 2020) or clustering data (Ghorbani et al., 2019), but they lack direct control over the outcomes and does not ensure the clarity of the concepts. One promising strategy is to utilize LLMs to provide comprehensible concept candidates. (Menon and Vondrick, 2022) use human language as an internal representation for visual recognition, and create an interpretable concept bottleneck for downstream tasks. By basing the decision on those comprehensible concepts, the model architecture itself is provided with better transparency. Similarly, a recent approach Labo (Yang et al., 2023a) constructs high-performance CBMs without manual concept annotations. This method controls the concept selection in bottlenecks by generating candidates from the LLMs, which contain significant world knowledge (Petroni et al., 2019) that can be explored by prompting a string prefix. Human studies further indicate that those LLM-sourced bottlenecks are much factual and groundable, maintaining great inherent interpretability for model designs. Besides the concept-based models, another promising strategy is to employ LLMs to enhance the conventional architectures that are inherently interpretable, such as GAMs and Decision Trees (DTs). (Singh et al., 2023) leverages the knowledge captured in LLMs to enhance GAMs and DTs, where LLMs are only involved during the augmented model training instead of the inference process. For GAMs training, LLMs can provide decoupled embeddings for enhancement. For DTs training, LLMs are able to help generate improved features for splitting. The LLM-augmented GAMs and DTs enable full transparency, where only the summing coefficients and input key phrases are required for interpretation. With the extra information from LLMs, augmented GAMs and DTs are capable of achieving better generalization performance compared with non-augmented ones.
10.2. Designing Interpretable AI Workflows with LLM Agents
@@
Traditional deep models are usually designed in an end-to-end manner. The internal workflows are not quite understandable to general users. By utilizing common-sense world knowledge, LLMs can break down complex problems into smaller ones and organize the workflows among them, leading to a more interpretable design of AI systems (Feng et al., 2023). A recent example of interpretable AI workflow design comes from  (Shen et al., 2024), where an LLM-powered agent leverages ChatGPT to integrate various off-the-shelf AI models (e.g., from Hugging Face (Jain, 2022)) to handle different downstream application tasks. In order to handle complicated tasks in a transparent workflow, LLMs play a pivotal role in coordinating external models with language mediums to harness their powers. By planning the target task, selecting candidate models, executing decomposed subtasks and summarizing responses, LLMs can help disassemble tasks based on user requests, and assign appropriate models to the tasks based on the model descriptions. Similarly, to transparentize the workflow, (Liu et al., 2023b) introduces a task decomposer to analyze the user prompts and break it down into a number of subtasks for solving using LLMs. Each subtask is well managed and attributed with description, domain, inputs, and outputs. In this way, the AI systems are then capable of handling intricate user prompts with a step-by-step understandable workflow. Under the prompting paradigm, (Khot et al., 2022) also employs LLMs to solve complex tasks by decomposition. Drawing inspiration from software libraries where the workflows are trackable, the decomposer and shared subtasks are designed in a modular manner. One step further, (Wang et al., 2024a) introduces an interactive planning approach for complex tasks, which enhances the error correction on initial LLM-generated plans by integrating plan execution descriptions and providing self-explanation of the feedback. Such an interactive nature enables better workflow transparency in long-term planning and multi-step reasoning task scenarios.
10.3. Challenges
10.3.1. Planning Feasibility in Complicated Scenarios
@@
Despite the task planning capability of LLMs, it is still challenging to apply to certain scenarios in real-world applications due to feasibility issues. One typical scenario is the few-shot planning cases (Guo et al., 2023), where acquiring large datasets for training is either impractical or cost-prohibitive, thus making feasible planning about unseen cases from sparse exemplars extremely challenging. To better assist the interpretable designs, LLM planning needs to generalize well without extensive supervision and is expected to have the ability to integrate information from prior experiences as well as knowledge. Besides, another important scenario lies in the dynamic planning settings (Dagan et al., 2023), in which LLMs integrate feedback from the environment iteratively, letting the agent take thinking steps or augment its context with a reasoning trace. Dynamic scenarios urgently and frequently involve high computational costs resulting from the iterated invocations of LLMs, and still face challenges in dealing with the limits of the context window and recovering from hallucinations on planning.
10.3.2. Assistance Reliability with Knowledge Gaps
@@
LLMs exhibit remarkable proficiency in encapsulating real-world knowledge within their parameters, but they resort to hallucinations and biases with high confidence when certain knowledge is missing or unreliable. As a result, a crucial research challenge keeps rising, i.e., how to effectively detect and mitigate the LLM knowledge gaps from humans when employing LLMs for designs. We will need further research on evaluating and developing robust LLM mechanisms to address the knowledge-gapping problems, with the goal of helping improve LLM reliability, reducing hallucinations and mitigating biases. Furthermore, the intersections between the knowledge gaps and the safety aspects are also a great challenge to be solved, which may pose some security concerns, especially when using LLMs for downstream models or workflow designs.
11. Emulating Humans with LLMs for XAI
@@
This section discusses how LLMs can play the role of humans in serving XAI development. Building explainable models requires two main steps where humans are in the loop: (1) collecting a dataset with human-annotated rationales to train the models; (2) collecting human feedback on the quality of explanations produced by the models for evaluation. The significant cost and time required for human involvement raise the main challenge in scaling up this procedure. LLMs emerge as a promising solution to this challenge, thanks to their capability to emulate human reasoning and produce responses that closely resemble human-generated content. In the following, we introduce the methods that demonstrate LLMs’ ability to generate human-like annotations and feedback for XAI.
11.1. Emulating Human Annotators for Training Explainable Models
@@
Incorporating human-understandable rationales into model development has shown its effectiveness in enhancing both the transparency and performance of the system for various NLP tasks, such as question answering (Li et al., 2018; Wu et al., 2020a), sentiment analysis (Du et al., 2019b; Antognini and Faltings, 2021), and common sense reasoning (Rajani et al., 2019; Camburu et al., 2021). We use the term rationales to describe supportive evidence that justifies the connection between inputs and outputs (Gurrapu et al., 2023). Traditionally, the rationales are collected by leveraging human annotations (Camburu et al., 2018; Wang et al., 2019) or applying expert-designed rules (Alhindi et al., 2018; Li et al., 2018), resulting in expensive costs or limited quality. Recently, researchers in automatic annotation (Ding et al., 2022; Belal et al., 2023; Gilardi et al., 2023) have begun to explore the potential of leveraging advanced LLMs to emulate human annotators in annotating the target labels of task-specific examples. These studies found that advanced LLMs show comparable annotation qualities against average crowd human annotators on most tasks with a lower cost, pointing out the scalability of using machine-emulated annotators. Inspired by these works, some studies (Huang et al., 2023b, a; Pang et al., 2024; Yang et al., 2024; Wu et al., 2024a; Zhang et al., 2024b) attempt to leverage advanced LLMs to collect rationales by applying the chain-of-thought prompting. Specifically, researchers provide several input-rationale-output demonstrations within the input text to prompt the LLMs to generate rationale and output for an unlabeled input instance. The quality of such annotated rationales largely relies on the in-context learning capabilities of LLMs, leading to uncontrollable annotation quality on uncommon tasks. Other scholars (Yao et al., 2023a; Chen et al., 2023b; Luo et al., 2023) propose a human-in-the-loop LLM-based annotation framework based on the active-learning architecture. This framework initially collects a small seed dataset with human-annotated rationales and labels. This seed dataset is used to train an explainable classifier for this downstream task. Then, each unlabeled sample is passed through the trained explainable classifier. This is followed by a selection strategy that chooses representative samples according to metrics such as explanation plausibility, prediction uncertainty, and sample diversity. Finally, LLMs are leveraged to annotate the rationales and labels of these selected unlabeled samples. This procedure could be repeated multiple times, and the trained explainable classifier from the latest time is the final output of this framework. Compared with other methods, this approach balances the annotation quality and the cost budget in developing explainable models by using LLM-emulated annotators.
11.2. Emulating Human Feedback for Evaluating Explainable Models
@@
The explanations generated by the explainable models could be classified into two categories: extractive and abstractive (Gurrapu et al., 2023). Extractive explanations derive directly from the input data, exemplified by attribution-based methods that emphasize specific segments of the input text. In contrast, abstractive explanations are generated in a free-form text manner, such as chain-of-thought (CoT) responses (Wei et al., 2022), offering a more nuanced interpretation. The quality of extractive explanations is typically assessed through their agreement with human-annotated rationales (DeYoung et al., 2020), such as accuracy, recall, and precision. However, evaluating abstractive explanations presents a significant challenge, as it is impractical to exhaustively evaluate all reasonable abstractive results comprehensively. To automatically assess abstractive explanations, early studies first collect some free-text rationales, and then compute the cosine similarity between embeddings of explanations and rationales (Cheng et al., 2023; Li et al., 2023d). A higher similarity between the abstraction explanation and the annotated rationales indicates a more transparent model. Recently, some researchers have followed the LLM-as-a-judge framework to ask LLM to judge the model explanations without referring to any human-annotated rationales (Miao et al., 2023; Bills et al., 2023; Zheng et al., 2023; Saha et al., 2025; Wei et al., 2024b), emphasizing the potential of emulating human feedback with advanced LLMs.
11.3. Challenges
11.3.1. Uncontrollable Credibility of Emulation